OpenStudy (anonymous):

A vector field F has constant strength <0,0,b>. Calculate the flux of F through the rectangle A=(6,0,4) B=(6,3,0) C=(0,3,0) and D=(0,0,4)

OpenStudy (anonymous):

I've got the parameterization of the rectangle to be G(x,y)=<x,y,4-4/3y> and the n to be n=<0,4/3,1>

OpenStudy (anonymous):

Solving the vector surface integral as \[\int\limits_{0}^{3}\int\limits_{0}^{6}<0,0,b>dotproduct<0,4/3,1>dxdy\] gives me an answer of 18b. However the answer I have in my book is b(pi). What am i doing wrong?